OpenStudy (anonymous):

( 2 + √6 ) (4 - √6) I NEED HELP WITH THIS QUESTION PLEASE , SOMEONE HELP

OpenStudy (gw2011):

(2+sq rt 6)(4-sq rt 6)=8-2(sq rt 6)+4(sq rt 6)-6 =2+2(sq rt 6) =2(1+sq rt 6)